# The Importance of Being Discrete: Measuring the Impact of Discretization in End-to-End Differentially Private Synthetic Data

### v2

@@ Abstract
In this paper, we present an extensive measurement study of four discretization strategies in the context of DP marginal generative models. More precisely, we design DP versions of three discretizers (uniform, quantile, and k-means) and reimplement the PrivTree algorithm. We find that optimizing both the choice of discretizer and bin count can improve utility, on average, by almost 30% across six DP marginal models, compared to the default strategy and number of bins, with PrivTree being the best-performing discretizer in the majority of cases. We demonstrate that, while DP generative models with non-private discretization remain vulnerable to membership inference attacks, applying DP during discretization effectively mitigates this risk. Finally, we propose an optimized approach for automatically selecting the optimal number of bins, achieving high utility while reducing both privacy budget consumption and computational overhead.
@@ 1 Introduction
Ostensibly, in the context of tabular data, DP synthetic data presents several benefits: high utility on downstream tasks, integration in existing data pipelines as a drop-in replacement for real data, ability to be safely regenerated/reused, etc. [36, 70, 17]. DP synthetic data is being adopted beyond academic research [36, 8, 33], by startups [18, 81], corporations [55, 34, 71], government agents [59, 60, 56, 63], regulators [83, 84, 17], and nonprofits [70, 85, 62].
While there is no single universally “best” model [36], benchmark studies [80, 21] have consistently ranked marginal models (including graphical [96, 51] and workload/query-based approaches [5, 42, 52]) as providing the best privacy-utility tradeoffs compared to deep-learning based techniques (e.g., Generative Adversarial Networks or Diffusion models). Marginal models are particularly effective for relatively small tabular datasets (fewer than 32 features) and simple tasks, e.g., capturing summary statistics and replicating them in the synthetic data [80, 21]. Broadly, they follow the select–measure–generate paradigm [51, 48]: i) select a collection of marginals of interest, ii) measure and add noise to the marginals, and iii) generate synthetic data that preserves them.
@@
Next, we measure the utility of the four discretization strategies on five controlled distributions with varying modality and skewness. We then assess their performance vis-à-vis five state-of-the-art marginal generative models (PrivBayes [96], MST [51], RAP [5], GEM [42], and AIM [52]), as well as a simple baseline model, dp-synthpop [45], on three popular datasets (Adult, Gas, and Wine). We experiment with a wide range of bin counts and evaluate four existing methods as well as a new optimized approach for automatically selecting the number of bins to use. We also show that the latter yields the best utility.
Finally, we examine the impact of applying DP to data domain extraction with different discretizers on the privacy leakage from the synthetic data using a well-known inference attack. Throughout our experiments, we evaluate the discretizers and generative models under varying privacy budgets – in total, we fit over 300,000 discretizers and train around 200,000 generative models.
@@
    Marginal models are highly sensitive to the discretization strategy and number of bins used. Their utility follows an inverted u-shaped trend as the number of bins increases: it initially improves but then degrades when too many bins introduce too much noise, as illustrated in Figure 1 (RQ2).
@@
    Optimizing both the discretizer and the number of bins can significantly help utility, with an improvement ranging from 9.28% to 43.54% compared to the default discretization (uniform with 20 bins), as shown in Figure 2 (RQ2).
@@
    Our optimized method for automatic bin count selection outperforms alternatives and provides a 41% normalized utility improvement over the default discretization (uniform with 20 bins) (RQ3).
@@
    Although extracting the data domain with DP slightly reduces utility ($\approx$4on average), it substantially reduces inference attacks’ success from 100% (perfect inference) to $\approx$50(near-random guessing) compared to non-private extraction (RQ4).
Overall, our study demonstrates the importance of discretization in end-to-end DP pipelines for synthetic tabular data generation, an essential step under-explored and un-optimized by prior work. By doing so, we not only show that substantial improvements in synthetic data utility can be made but also that discretization itself needs to be done in a DP manner, or these pipelines could be left exposed to serious privacy vulnerabilities. We are confident that our work will help developers and practitioners be more conscious of the implementation details surrounding DP synthetic tabular data generation and result in models with better privacy-utility tradeoffs.
@@ 2 Background
Discretization. The process of mapping a continuous value $\underline{x}\leq x\leq\overline{x}$ into a discrete value $\hat{x}\in\{1,2,\dots,b\}$ is known as discretization. This involves partitioning the domain of $x$ ($[\underline{x},\overline{x}]$) into $b$ non-overlapping intervals, or bins, defined by their edges $\{e_{1},e_{2},\dots,e_{b+1}\}$, where $e_{1}=\underline{x}$ and $e_{b+1}=\overline{x}$ denote the minimum and maximum bounds of the domain, respectively. The bin edges satisfy the properties $e_{1}<e_{2}<\dots<e_{b+1}$, ensuring full coverage of the range, i.e., $\bigcup_{i=1}^{b}[e_{i},e_{i+1}]=[e_{1},e_{b+1}]$. A continuous value $x$ is mapped to $i$ if $x\in[e_{i},e_{i+1})$ $\forall i\neq b$ and $\hat{x}=b$ if $x\in[e_{b},e_{b+1}]$. The discretized value can also be mapped back onto the continuous range, typically by sampling from the corresponding interval.
@@
Privacy Attacks. We use membership inference attacks (MIAs) [74, 30] to estimate the privacy leakage from a trained generative model. In an MIA, an adversary attempts to determine whether a specific target record was included in the training dataset of the generative model (i.e., $\mathbf{x}_{T}\in D_{train}$), which closely aligns with the DP definition.
The attack can be formulated as a distinguishing game, where the adversary is presented with either $D_{synth}$, generated by a model trained on $D_{train}$, or $D_{synth}^{\prime}$, generated by a model trained on $D_{train}^{\prime}=D_{train}\setminus\mathbf{x}_{T}$, and has to predict which training dataset was used. The adversary’s predictions, $Pred=\{pred_{1},pred_{2},\dots\}$ and $Pred^{\prime}=\{pred^{\prime}_{1},pred^{\prime}_{2},\dots\}$, are analyzed, and their performance is quantified using the AUC score, which measures their success in distinguishing between the two scenarios.
@@ 3 Evaluation Framework
3.1 Discretization
We experiment with four different discretization strategies, two sampling techniques, and four methods for determining the optimal number of bins. The DP adaptations (of most) of these methods, which have primarily been studied in non-DP contexts, are introduced later in Section 4.
Discretizers. They work independently on a single continuous column (one-dimension) of the dataset, taking as input the values of the columns, their domains/ranges, and the desired number of bins ($b$). We experiment with the following four discretizers:
  • $\bullet$
    uniform splits the data domain into $b$ intervals of equal width.
  • $\bullet$
  • $\bullet$
  • $\bullet$
@@ 3.1 Discretization
    Uniform samples uniformly within the range defined by the bin edges.
@@
Optimal Number of Bins. In Table 1, we list various methods for selecting the optimal bin count. These typically consider various aspects of the data distribution, such as the number of records, moments, range, etc.
@@ 3.2 DP Generative Models
RAP [5] iteratively identifies high error queries, projects them into a continuous relaxation of the data domain (an extension of the projection mechanism from [58]), and optimizes them through gradient-based methods. Each round involves private selection (via a tailored Report Noisy Max mechanism [16] with Gumbel noise) and measurements via a Gaussian mechanism. Finally, randomized rounding is applied to revert the generated data to the original domain.
GEM [42] follows a similar iterative approach but uses a neural network, modeled after a GAN’s generator [27], to implicitly learn a product distribution representation over the attributes in the data domain. This learned distribution can then directly be used to respond to marginal queries. To satisfy DP, it employs the Exponential mechanism for private sampling and the Gaussian mechanism for noisy measurements.
AIM [52] similar to MST, uses the Private-PGM algorithm to generate synthetic data that preserves noisy measurements. However, AIM differs by dynamically selecting marginals of interest based on their overall importance, in contrast to the static selection method used by MST. The process of selecting and measuring the relevant marginals (with the Gaussian mechanism) is then repeated iteratively until the privacy budget is exhausted or a stopping condition is met.
@@ 3.3 Metrics
Record Similarity (RS). RS is used when comparing two sets of records with a one-to-one correspondence. It is computed as the average normalized L1 distance, i.e.,
@@ 3.4 Datasets
Controlled Distributions. We create five datasets, namely, monotone, normal, beta, skewed, mixture, and imbalanced, reported in Table 2. (Also see Section 5.1 for visual details.) We also include an additional dataset, uniform, solely for visualization purposes. The datasets are clipped and rescaled to fit within the $[-10,10]$ domain, and include a varying number of records, i.e., 500, 1k, 10k, and 100k.
Real Datasets. We use common real datasets, i.e., Adult, Gas, and Wine from UCI’s ML Repository [13], all of them having associated binary classification tasks – see Table 3.
@@
3.5 Experimental Settings
We introduce four experimental settings, three focused on utility and one on privacy, as visualized in Figure 3. In all utility settings, we fit three generative models, generate three synthetic datasets per trained model, and report aggregated metrics based on the generated 9 synthetic datasets.
Capturing Controlled Distributions (US1). We measure how much utility DP discretizers preserve in controlled settings (cf. Section 5.1). The goal is to isolate their performance under controlled conditions. To do so, we discretize the input datasets using the four discretizers and sample uniformly from each bin, assuming a known data domain. The entire privacy budget is spent in the discretization step and utility is measured by averaging all applicable one-dimensional metrics – record similarity, maximum percentile distance, discriminator similarity, and 1d query similarity. We use the five controlled distributions and their size variations.
Modeling Controlled Distributions (US2). Next, we evaluate the utility retained when baseline DP generative modeling is applied alongside DP discretization in controlled settings (as explored in Section 5.2 and 5.3). This simulates the simplest generative modeling scenario for one-dimensional data. The datasets are discretized using the four discretizers, and the distributions are modeled using the Laplace mechanism, as done by dp-synthpop and PrivBayes when applied to one-dimensional data. The privacy budget is split, with 10% allocated to discretization and 90% to modeling. Unless stated otherwise, we assume the data domain is known (i.e., not extracted from the data) and sample uniformly from each bin. Utility is measured using the same metrics as in US1, except for record similarity. We use the same five controlled distributions as in US1.
@@ 3.5 Experimental Settings
Modeling Real Datasets (US3). We also assess the utility of DP generative models combined with DP discretization on real datasets (used in Section 5.25.3, and 5.4) to simulate realistic end-to-end DP generative model deployments. The datasets are preprocessed using the four discretizers and used with six DP generative models. As in previous work, we apply the models with their default hyperparameters [80, 21]. The privacy budget is distributed 10% to discretization, divided equally across columns, and 90% to modeling. Unless otherwise specified, the data domain is assumed to be known (we use DP extraction in Section 5.4), and uniform sampling is used from the bins (we compare uniform and mixture sampling in Section 5.2). We measure utility with all listed metrics (without record similarity). For this setting, we use the three real datasets: Adult, Gas, and Wine.
@@
Privacy Leakage (PS1). In the privacy setting, we evaluate the domain’s contribution to overall privacy leakage within the end-to-end DP pipeline using MIAs (cf. Section 5.4). We assess three domain extraction strategies: assuming a given domain, extracting it directly from the input data, or extracting it with a DP. Due to computational constraints, we focus on four discretizers combined with two generative models, i.e., PrivBayes and MST, whose privacy properties have been extensively studied [4, 26].
@@
We use the GroundHog membership inference attack [76], which is arguably the most popular attack targeting synthetic data generation [32, 4, 22]. First, we identify a vulnerable record as the target, selecting the data point furthest from all others in the training set [54] and ensuring it lies outside their domain. Next, we train two sets of shadow models (200 models each): one trained on the full dataset, including the target record, and the other excluding it. The discretization privacy budget is allocated based on the domain extraction strategy: 100% to discretization when the domain is provided or extracted non-privately, and split 50/50 between domain extraction and discretization when the domain is extracted with DP. We then generate synthetic datasets. We extract statistical features (minimum, maximum, mean, median, and standard deviation) from each column of the synthetic datasets. Half of these features are used to train a classifier, and the AUC score is evaluated and reported on the remaining data. For this evaluation, we use a single dataset, Wine.
@@ 4 DP Discretization in End-to-End DP Generative Pipeline
In this section, we introduce the Differentially Private (DP) discretization components used in place of the non-DP versions introduced in Section 3.1, which ultimately allows for the construction of an end-to-end DP generative pipeline. Specifically, we present DP variants for domain extraction, three discretizers (uniform, quantile, k-means), two sampling methods (uniform and mixture), and a strategy for selecting the optimal number of bins. Throughout, we use primitives of two well-known open-source libraries, namely, Harvard’s OpenDP [64] and IBM’s Diffprivlib [31].
DP Domain Extraction. We implement the algorithm by Desfontaines [9] to estimate the domain of numerical data given a privacy budget $\epsilon$. It derives bounds using a noisy histogram over an exponential range $[-2^{m},2^{m}]$, with $m$ typically set to 32. The bounds are determined by iteratively reducing a threshold until at least one bin exceeds it, using the highest and lowest bin edges above the threshold serving as the domain bounds. [^1] Inspired by OpenDP, see https://github.com/opendp/smartnoise-sdk/blob/main/sql/snsql/sql/˙mechanisms/approx˙bounds.py. Note that this method may fail if $\epsilon$ is too small; however, since this does not affect the conclusions drawn from our measurement study, addressing this potential issue is left for future work.
@@
  • DP-uniform does not consume any privacy budget and determines bin edges only relying on the provided data domain.
@@
  • PrivTree [95] is inherently DP. The threshold parameter $\tau$ is set to ${1}/{b}$, making $b$ an upper limit for the actual number of bins produced.
@@
  • DP-Uniform sampling does not use $\epsilon$ as it samples uniformly from the corresponding bin edges.
@@
DP Selection of the Optimal Number of Bins. Finally, we optimize the Rice rule’s formula ($b^{\mathbf{RL}}$) in the context of DP, as the alternatives ($b^{\mathbf{DF}}$, $b^{\mathbf{FDR}}$, and $b^{\mathbf{SHI}}$) depend on aspects of the data distribution, requiring additional privacy budget expenditure. We incorporate a correction term, taking into account the available privacy budget $\epsilon$, and define $b^{\mathbf{RL-OPT}}:=\dfrac{2\sqrt[3]{n}}{1+exp(-\epsilon)}$. The correction term aims to reduce the number of bins for small $\epsilon$.

### v4

@@ Abstract
In this paper, we present an extensive measurement study of four discretization strategies in the context of DP marginal generative models. More precisely, we design DP versions of three discretizers (uniform, quantile, and k-means) and reimplement the PrivTree algorithm. We find that optimizing both the choice of discretizer and bin count can improve utility, on average, by almost 30% across six DP marginal models, compared to the default strategy and number of bins, with PrivTree being the best-performing discretizer in the majority of cases. We demonstrate that, while DP generative models with non-private discretization remain vulnerable to membership inference attacks, applying DP during discretization effectively mitigates this risk. Finally, we improve on an existing approach for automatically selecting the optimal number of bins, and achieve high utility while reducing both privacy budget consumption and computational overhead.
footnotetext: Published in the Proceedings of the 32nd ACM Conference on Computer and Communications Security (ACM CCS 2025).
@@ 1 Introduction
In the context of tabular data, DP synthetic data ostensibly offers several benefits: high utility on downstream tasks, seamless integration into existing data pipelines, ability to be safely regenerated/reused, etc. [37, 71, 17]. DP synthetic data is being adopted beyond academic research [37, 8, 34], by startups [18, 82], corporations [56, 35, 72], government agents [60, 61, 57, 64], regulators [84, 85, 17], and nonprofits [71, 86, 63].
While there is no single universally “best” model [37], benchmark studies [81, 21], real-world deployments [64, 85, 31], and NIST competitions [60, 61] have consistently ranked marginal models (including graphical [97, 52] and workload/query-based approaches [5, 43, 53]) as providing the best privacy-utility tradeoffs compared to deep-learning based techniques. Marginal models are particularly effective for relatively small tabular datasets (fewer than 32 features) and simple tasks, e.g., capturing summary statistics and replicating them in the synthetic data [81, 21]. Broadly, they follow the select–measure–generate paradigm [52, 49]: i) select a collection of marginals of interest, ii) measure and add noise to the marginals, and iii) generate synthetic data that preserves them.
@@
Next, we measure the utility of the four discretization strategies on five controlled distributions with varying modality and skewness. We then assess their performance vis-à-vis five state-of-the-art marginal generative models (PrivBayes [97], MST [52], RAP [5], GEM [43], and AIM [53]), as well as a simple baseline model, dp-synthpop [46], on three popular datasets (Adult, Gas, and Wine). We experiment with a wide range of bin counts and evaluate four existing methods as well as an optimized approach for automatically selecting the number of bins to use. We also show that the latter yields the best utility.
Finally, we examine the impact of applying DP to data domain extraction with different discretizers on the privacy leakage from the synthetic data using two well-known inference attacks. Throughout our experiments, we evaluate the discretizers and generative models under varying privacy budgets – in total, we fit over 300,000 discretizers and train around 200,000 generative models.
@@
    Marginal models are highly sensitive to the discretization strategy and number of bins used. Their utility follows an inverted u-shaped trend as the number of bins increases: it initially improves but then degrades when too many bins introduce too much noise, as illustrated in Figure˜2 (RQ2).
@@
    Optimizing both the discretizer and the number of bins can significantly help utility, with an improvement ranging from 9.28% to 43.54% compared to the default discretization (uniform with 20 bins), as shown in Figure˜2 (RQ2).
@@
    Our optimized method for automatic bin count selection outperforms alternatives and provides a 41% normalized utility improvement relative to the range between the default and the best achievable utility via exhaustive search (RQ3).
@@
    Although extracting the data domain with DP slightly reduces utility ($\approx$4% on average), it substantially reduces inference attacks’ success from 100% (perfect inference) to $\approx$50% (near-random guessing) compared to non-private extraction (RQ4).
Overall, our study demonstrates the importance of discretization in end-to-end DP pipelines for synthetic tabular data generation, an essential step under-explored and un-optimized by prior work. By doing so, we not only show that substantial improvements in synthetic data utility can be made but also that discretization itself needs to be done in a DP manner, or these pipelines could be left exposed to serious privacy vulnerabilities. We are confident that our work will help developers and practitioners be more conscious of the implementation details surrounding DP synthetic data generation and result in models with better privacy-utility tradeoffs.
@@ 2 Background
[…]
@@
Privacy Attacks. We use membership inference attacks (MIAs) [75, 30] to estimate the privacy leakage from a trained generative model. In an MIA, an adversary attempts to determine whether a specific target record was included in the training dataset of the generative model (i.e., $\mathbf{x}_{T}\in D_{train}$), which closely aligns with the DP definition. The attack can be formulated as a distinguishing game, where the adversary is presented with either $D_{synth}$, generated by a model trained on $D_{train}$, or $D_{synth}^{\prime}$, generated by a model trained on $D_{train}^{\prime}=D_{train}\setminus\mathbf{x}_{T}$, and has to predict which training dataset was used. The adversary’s predictions, $Pred=\{pred_{1},pred_{2},\dots\}$ and $Pred^{\prime}=\{pred^{\prime}_{1},pred^{\prime}_{2},\dots\}$, are analyzed, and their performance is quantified using the AUC score, which measures their success in distinguishing between the two scenarios.
@@ 3 Evaluation Framework
3.1 Discretization
We experiment with four different discretization strategies, two sampling techniques, and four methods for determining the optimal number of bins. The DP adaptations (of most) of these methods, which have primarily been studied in non-DP contexts, are introduced later in Section˜4.
Discretizers. The discretizers work independently on a single continuous column (one-dimension) of the dataset, taking as input the column values, the domain/range of the column, and the desired number of bins ($b$). We evaluate the following four discretizers:
  • $\bullet$
    uniform splits the data domain into $b$ intervals of equal width.
  • $\bullet$
  • $\bullet$
  • $\bullet$
@@ 3.1 Discretization
    Uniform samples uniformly within the defined bin edges.
@@
Optimal Number of Bins. In Table˜1, we list various methods for selecting the optimal bin count. These typically consider various data distribution aspects, e.g., number of records, moments, range.
@@ 3.2 DP Generative Models
We select the five DP marginal models, as they have been published in top-tier venues, are among the most widely used/cited, and have open-source implementations. Overall, our experiments are representative of the space, and we are not knowingly leaving out any alternatives that might yield significantly different results.
@@
RAP [5] iteratively identifies high error queries, projects them into a continuous relaxation of the data domain (an extension of the projection mechanism from [59]), and optimizes them through gradient-based methods. Each round involves private selection (via a tailored Report Noisy Max mechanism [16] with Gumbel noise) and measurements via a Gaussian mechanism. Finally, randomized rounding is used to map the generated data to the original domain.
GEM [43] follows a similar iterative approach but uses a neural network, modeled after a GAN’s generator [27], to implicitly learn a product distribution representation over the attributes in the data domain. The learned distribution can directly be used to respond to marginal queries. To satisfy DP, it employs the Exponential mechanism for sampling and the Gaussian mechanism for measurements.
AIM [53] similar to MST, uses the Private-PGM algorithm to generate synthetic data that preserves noisy measurements. However, AIM dynamically selects marginals of interest based on their overall importance, in contrast to the static selection method used by MST. The process of selecting and measuring the relevant marginals (with the Gaussian mechanism) is then repeated iteratively until the privacy budget is exhausted or a stopping condition is met.
@@ 3.3 Metrics
Record Similarity (RS). RS is used when comparing two sets of records with a one-to-one correspondence. It is computed as the average normalized L1 distance:
@@ 3.4 Datasets
Controlled Distributions. We create five datasets, namely, monotone, normal, beta, skewed, mixture, and imbalanced, reported in Table˜2. (Also see Section˜5.1 for visual details.) We also include an additional dataset, uniform, solely for visualization purposes. The datasets are clipped and rescaled to fit within the [10,10][-10,10] domain, and include a varying number of records, i.e., 500, 1k, 10k, and 100k.
@@
Real Datasets. We use common real datasets, i.e., Adult, Gas, and Wine from UCI’s ML Repository [13], all of them having associated binary classification tasks – see Table˜3.
3.5 Experimental Settings
We introduce four experimental settings, three focused on utility and one on privacy, as visualized in Figure˜3. In all utility settings, we fit three generative models, generate three synthetic datasets per trained model, and report aggregated metrics based on the generated 9 synthetic datasets. We average across the metrics to provide a concise unbiased score, which covers various use cases and enables straightforward comparison across discretizers/models.
Capturing Controlled Distributions (US1). We measure how much utility DP discretizers preserve in controlled settings (cf. Section˜5.1). The goal is to isolate their performance under controlled conditions. To do so, we discretize the input datasets using the four discretizers and sample uniformly from each bin, assuming a known data domain. The entire privacy budget is spent in the discretization step and utility is measured by averaging all applicable one-dimensional metrics – record similarity, maximum percentile distance, discriminator similarity, and 1d query similarity. We use the five controlled distributions and their size variations.
Modeling Controlled Distributions (US2). Next, we evaluate the utility retained when baseline DP generative modeling is applied alongside DP discretization in controlled settings (as explored in Section˜5.2 and 5.3). This simulates the simplest generative modeling scenario for one-dimensional data. The datasets are discretized using the four discretizers, and the distributions are modeled using the Laplace mechanism, as done by dp-synthpop and PrivBayes when applied to one-dimensional data. The privacy budget is split, with 10% allocated to discretization and 90% to modeling. Unless stated otherwise, we assume the data domain is known (i.e., not extracted from the data) and sample uniformly from each bin. Utility is measured using the same metrics as in US1, except for record similarity. We use the same five controlled distributions as in US1.
@@ 3.5 Experimental Settings
Modeling Real Datasets (US3). We also assess the utility of DP generative models combined with DP discretization on real datasets (used in Section˜5.25.3, and 5.4) to simulate realistic end-to-end DP generative model deployments and test whether our findings in US1/US2 generalize. The datasets are preprocessed using the four discretizers and used with six DP generative models. As in previous work, we apply the models with their default hyperparameters [81, 21]. The privacy budget is distributed 10% to discretization, divided equally across columns, and 90% to modeling. Unless otherwise specified, the data domain is assumed to be known (we use DP extraction in Section˜5.4), and uniform bin sampling is used (we compare uniform and mixture sampling in Section˜5.2). We measure utility with all listed metrics (without record similarity). For this setting, we use the three real datasets: Adult, Gas, and Wine.
@@
Privacy Leakage (PS1). In the privacy setting, we evaluate the domain’s contribution to overall privacy leakage within the end-to-end DP pipeline using MIAs (cf. Section˜5.4). We assess three domain extraction strategies: assuming a given domain, extracting it directly from the input data, or extracting it with a DP. Due to computational constraints, we focus on four discretizers combined with two generative models, i.e., PrivBayes and MST, whose privacy properties have been extensively studied [4, 26].
@@
We use the GroundHog [77] and Querybased [33] membership inference attacks, which are arguably the most popular attack targeting synthetic data generation [4, 22]. First, we identify a vulnerable record as the target, selecting the data point furthest from all others in the training set [55] and ensuring it lies outside their domain. Next, we train two sets of shadow models (200 models each): one trained on the full dataset, including the target record, and the other excluding it. The discretization privacy budget is allocated based on the domain extraction strategy: 100% to discretization when the domain is provided or extracted non-privately, and split 50/50 between domain extraction and discretization when the domain is extracted with DP. We then generate synthetic datasets. For GroundHog, we extract statistical features (minimum, maximum, mean, median, and standard deviation) from each column of the synthetic data, while for Querybased, we extract the counts of all possible subsets of columns of the synthetic data that match the corresponding values in the target record. Half of these features are used to train a classifier, and the AUC score is evaluated and reported on the remaining data. We use a single dataset, Wine.
@@ 4 DP Discretization in End-to-End DP Generative Pipeline
In this section, we introduce the Differentially Private (DP) discretization components used in place of the non-DP versions introduced in Section˜3.1, which ultimately allows for the construction of an end-to-end DP generative pipeline. Specifically, we present DP variants for domain extraction, three discretizers (uniform, quantile, k-means), two sampling methods (uniform and mixture), and a strategy for selecting the optimal number of bins. Throughout, we use primitives of two well-known open-source libraries, namely, Harvard’s OpenDP [65] and IBM’s Diffprivlib [32].
DP Domain Extraction. We implement the algorithm by Desfontaines [9] to estimate the domain of numerical data given a privacy budget $\epsilon$. It derives bounds using a noisy histogram over an exponential range $2^{m},2^{m}]$, with $m$ typically set to 32. The bounds are determined by iteratively reducing a threshold until at least one bin exceeds it, using the highest and lowest bin edges above the threshold serving as the domain bounds. [^1] Inspired by OpenDP, see https://github.com/opendp/smartnoise-sdk/blob/main/sql/snsql/sql/˙mechanisms/approx˙bounds.py. Note that this method may fail if $\epsilon$ is too small; however, since this does not affect the conclusions drawn from our measurement study, addressing this potential issue is left for future work.
@@
  • (DP)-uniform does not consume any privacy budget and determines bin edges only relying on the provided data domain.
@@
  • PrivTree [96] is inherently DP. The threshold parameter $\tau$ is set to 1/b{1}/{b}, making $b$ an upper limit for the actual number of bins produced.
@@
  • (DP)-Uniform sampling does not use $\epsilon$ as it samples uniformly from the corresponding bin edges.
@@
DP Selection of the Optimal Number of Bins. Finally, we optimize the Rice rule’s formula ($b^{\mathbf{RL}}$) in the context of DP, as the alternatives ($b^{\mathbf{DF}}$, $b^{\mathbf{FDR}}$, and $b^{\mathbf{SHI}}$) depend on aspects of the data distribution, requiring additional privacy budget expenditure. We incorporate a correction term, taking into account the available privacy budget $\epsilon$, and define $b^{\mathbf{RL-OPT}}:=\frac{2\sqrt[3]{n}}{1+exp(-\epsilon)}$. The correction term aims to reduce the number of bins for small $\epsilon$.